### v2

# Non-equilibrium classical recombination in the expanding ultracold plasmas

Abstract
The efficiency of recombination is of crucial importance for the existence of ultracold plasmas (UCP), particularly, the ones formed in the magneto–optical traps. Unfortunately, the equilibrium thermodynamic treatment of the ionization–recombination processes is inappropriate for the evolving UCP clouds, while the straightforward kinetic simulation encounters the problem of huge difference in the spatial and temporal scales for free and bound motion of the electrons. As a result, only the “virtual” electron–ion pairs are usually reproduced in such modeling, and it is necessary to employ some heuristic criteria to identify them with the recombined atoms. It is the aim of this paper to present the first successful ab initio simulation of the non-equilibrium recombination in the evolving UCP plasmas. We employed a special algorithm, which is based on using the “scalable” reference frame, co-moving with the expanding substance. Then, the recombination events are identified by a series of sharp equidistant peaks in the kinetic and/or potential energies, which are caused by the captured electrons passing near the pericenters of their orbits; and this is confirmed by a detailed inspection of their trajectories. Thereby, we were able to trace the real—rather than “virtual”—electron–ion pairs, and the total efficiency of their formation was found to be about 20%, which is in agreement with the laboratory measurements.
I Introduction
The ultracold plasmas (UCP), formed after release of the bunches of photoionized atoms from the magneto-optical traps Killian et al. (1999), represent a relatively new branch of plasma physics, emerged in the very late 1990s and early 2000s Gould and Eyler (2001); Bergeson and Killian (2003); Killian (2007); Killian et al. (2007). The possibility of existence of such plasmas for an appreciable time interval crucially depends on the efficiency of recombination, since it might be intuitively expected that the rarefied charged particles with a sufficiently strong Coulomb coupling will quickly collapse into the neutral pairs.
It is interesting to mention that the first molecular-dynamic simulations of such plasmas started already in the early 1990s, i.e., a decade before the first experiments with UCP. Surprisingly, these simulations demonstrated that the ultracold plasmas experienced a very small recombination, i.e., could survive for a sufficiently long time Mayorov, Tkachev, and Yakovlenko (1994). Unfortunately, it was suspected by some opponents that the anomalous resistance of charged particles with respect to the recombination might be just an artifact of the particular numerical methods. For example, the recombined neural pairs could be efficiently destroyed by the collisions with walls of the computational box Ignatov et al. (1995); Mayorov, Tkachev, and Yakovlenko (1995), if the reflective boundary conditions were imposed there. Moreover, the anomalously small recombination rate might be caused just by the accumulation of numerical errors, preventing formation of the bound electron–ion pairs. This problem is especially severe because of the huge difference in the characteristic temporal and spatial scales for the free and bound motion of the charged particles in plasmas.
In view of the above-mentioned problem, the most of works on recombination in UCP employed the artificial two-step procedure: Firstly, the electron distribution function was simulated numerically by the molecular-dynamic approach, and then it was used for analytical estimates of the recombination rate, e.g., employing the idea of electron diffusion in the energetic space Bobrov et al. (2011).
On the other hand, as regards ab initio modeling of recombination, the most of works employed some additional criteria for the formation of the bound electron–ion pairs: For example, it was assumed that the act of recombination took place if an electron performed a certain number of revolutions about the nearest ion (e.g., 4 complete revolutions) Lankin and Norman (2009a). Yet another approach was to assume that all electrons located closer than a specified distance from the nearest ions (e.g., 20% of the Wigner–Seitz radius) will ultimately recombine Niffenegger, Gilmore, and Robicheaux (2011). These artificial criteria were necessary because the commonly-used numerical schemes did not allow ones to get sufficiently stable electron–ion pairs: they were destroyed rather quickly due to accumulation of the computational errors. Besides, these works usually performed the simulations within a box of a fixed size, which is poorly relevant to the actual experimental conditions, since the UCP clouds typically expand by one or two orders of magnitude in the course of experiments.
It is the aim of the present work to employ a specially designed numerical scheme that is able, firstly, to take into account the considerable degree of expansion of the plasma cloud and, secondly, to trace formation of the stable electron–ion pairs (which should be subsequently converted to neutral atoms due to the quantum processes). In fact, the basic idea of this algorithm was already used in our earlier paper Dumin (2011), which was aimed at ab initio simulation of the electron temperature evolution in UCP bunches released from the magneto-optical traps. The corresponding functional dependence obtained in that work, $T_{e}\propto t^{\alpha}$ with exponent $\alpha\in[-1.25,-1.08]$, was in perfect agreement with the experiment Fletcher, Zhang, and Rolston (2007), where $\alpha=-(1.2{\pm}0.1)$. It is important to emphasize that this coincidence was achieved exactly from “the first principles”, without postulation of any additional heat sources or sinks. As will be shown below, the same numerical scheme provides very good results also in the simulation of recombination.
II The algorithm of simulation
II.1 Geometrical configuration
Figure 1: Sketch of the plasma cloud (gray segment with a blue boundary) composed of the identical cells of size $L$, uniformly expanding with velocity $u_{0}$.
Following the design of typical UCP experiments, we shall consider a large volume of plasma expanding uniformly in all directions, starting from the center $r=0$. For the sake of simplicity, we assume that the cloud is uniform and has a sharp boundary at $r=R(t)$; see Fig. 1. Next, let this cloud be composed of the identical (or “mirror”) cubic cells filled with some number of ions $N$ and the same number of electrons, and $N_{\rm m}$ be the number of cells from the center to the boundary. Each cell expands with velocity $u_{0}$ with respect to its opposite face. So, at the cloud boundary $R=N_{\rm m}L$, the velocity becomes $V=N_{\rm m}\,u_{0}$. In fact, this scheme closely resembles the idea of Hubble-Lemaître expansion, commonly used in cosmology Zel’dovich and Novikov (1983).
It is well known that the initial thermal energy of plasma (which, in the case of photoionization, is concentrated mostly in electrons) quickly transforms into kinetic energy of macroscopic motion of the cloud; e.g., papers Vikhrov et al. (2020, 2021) and references therein. In other words, $m_{i}\,V^{2}\sim\,k_{\rm B}\,T_{e0}$, where $T_{e0}$ is the initial electron temperature, $V$ is the velocity of the cloud boundary, $m_{i}$ is the ionic mass, and $k_{\rm B}$ is Boltzmann constant. Consequently,
$V\sim\sqrt{k_{\rm B}\,T_{e0}/m_{i}}$ (1)
(we did not write here any numerical coefficient, since it depends substantially on the distribution of density inside a more realistic cloud). Just this formula can be used to link the experimental quantities with the simulation parameters.
II.2 Equations of motion
Figure 2: Sketch of calculation of the Coulomb sums, where each electron $e$ interacts not only with a specified ion $i$ inside the basic cell (solid arrow) but also with infinite number of its “mirror” images $i^{\prime}$ (dashed arrows).
The simulations are based on solving the classical equations of motion for $N$ electrons inside the “basic” cell:
$m_{e}\,\frac{d^{2}{\rm\bf r}_{i}}{dt^{2}}=\,{\rm\bf F}_{i}\,$, (2)
where ${\rm\bf r}_{i}$ are the electron coordinates, and $m_{e}$ is the electron mass. The Coulomb force ${\rm\bf F}_{i}$, acting on the $i$’th electron, is calculated taking into account all particles in the system i.e., not only within the same basic cell but also the all its “mirror” images (Fig. 2):
${\bf F}_{i}$ $=$ $\sum_{j=1}^{N}e^{2}\,\frac{{\bf R}_{j}-{\bf r}_{i}}{|{\bf R}_{j}-{\bf r}_{i}|^{3}}\;+\,\sum_{j=1,\,j\neq i}^{N}e^{2}\,\frac{{\bf r}_{i}-{\bf r}_{j}}{|{\bf r}_{i}-{\bf r}_{j}|^{3}}$ (3)
$+$ $\sum_{{\bf n}\,\in\,\mathbb{Z}^{3}\backslash\{{\bf 0}\}}\Bigg\{\sum_{j=1}^{N}e^{2}\frac{({\bf R}_{j}+L\,{\bf n})-{\bf r}_{i})}{\big|({\bf R}_{j}+L\,{\bf n})-{\bf r}_{i}\big|^{3}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\>+\;\sum_{j=1}^{N}\,e^{2}\frac{{\bf r}_{i}-({\bf r}_{j}+L\,{\bf n})}{\big|{\bf r}_{i}-({\bf r}_{j}+L\,{\bf n})\big|^{3}}\Bigg\}\>.$
[…]
Figure 3: Sketch of summation over the mirror cells (dotted squares) around the basic cell (solid square) in calculation of the Coulomb forces.
In fact, summation in formula (3) is performed over a finite number of mirror cells until the specified accuracy of convergence of the Coulomb forces is achieved. These mirror cells are arranged into the cubic shells of increasing radii around the basic cell, as illustrated in Fig. 3. Strictly speaking, the sums like (3) are conditionally convergent, i.e., the result can depend on the summation order. The method of summation depicted in the figure seems most reasonable from the physical point of view. Really, effect of the mirror cells located at the larger radii should be obscured by some additional physical perturbations, thereby providing a “physical regularization” of the conditionally-convergent Coulomb sums.
Although we integrated the equations of motion for a relatively small number of charged particles in the basic cell (e.g., a few dozens or a few hundreds), the total number of particles taken into account in calculation of the Coulomb sums was huge. For example, if the relative accuracy of calculations was specified to be about $10^{-4}$ or $10^{-5}$, then even at $N=10$ the total number of particles involved in the computation turned out to be from a few hundred thousand to a few million. Let us emphasize that this number was not fixed but chosen “dynamically” at each step of integration.
In fact, the same type of summation as in formula (3) is commonly implemented in the condensed-matter physics by the so-called Ewald method. (Its original idea is outlined, e.g., in textbook Ziman (1972), and a more elaborated version can be found in paper Demyanov and Levashov (2022).) Most probably, the same approach can be applied also to the problem under consideration, but this issue requires a further study.
II.3 The scalable reference frame
Yet another serious problem in simulation of the expanding plasma clouds is a considerable change in the spatial scale of the system—and, therefore, in the Coulomb forces—in the course of time, which makes it difficult to keep the accuracy of integration at a stable level. To get around this obstacle, we introduce the scalable reference frame, expanding with the average expansion rate of the plasma. Namely, length of the basic cell is assumed to increase as
$L(t)=L_{0}+u_{0}\,t,$ (4)
where $L_{0}$ is its initial value, and $u_{0}$ is the velocity of expansion. From the physical point of view, such a linear law corresponds to the inertial stage of the plasma motion, when the most part of its initial thermal energy was transformed into kinetic energy of the macroscopic expansion. In fact, this law is established very quickly, and it is commonly used in the interpretation of all experiments with UCP.
Besides, it is convenient to introduce the following dimensionless variables. Let the unit of length be the time-dependent characteristic interparticle separation:
$\tilde{l}(t)=L(t)/{(2N)^{1/3}}\,.$ (5)
The unit of time is taken to be proportional to the characteristic period of oscillations of an electron with mass $m_{e}$ in the field of the nearest ion at the initial instant of time:
$\tau=\big(m_{e}^{1/2}\!/e\big)\,{\tilde{l_{0}}}^{\,3/2}$ (6)
(as distinct from the unit of length, it is independent of time). From here on, all physical quantities normalized to $\tilde{l}$ and $\tau$ will be marked by asterisks.
As can be easily seen, the unit of length varies with the dimensionless time as
$\tilde{l}=\>\tilde{l_{0}}\,(1+\,u_{0}^{*}\,t^{*})\,$, (7)
where
$u_{0}^{*}=\,u_{0}\,\tau/L_{0}$ (8)
is the dimensionless velocity of expansion; while the dimensionless size of the basic cell remains constant:
$L^{*}=(2N)^{1/3}.$ (9)
In other words, the simulated particles perform their motion in a box of the fixed size (9), thereby mitigating the problem of large variation of the spatial coordinates in the case of considerable plasma expansion.
As usual, the periodic boundary conditions are imposed at the box boundaries: when one of the particles leaves the box through one of its faces, the identical particle with the same velocity enters this box through the opposite face.
The equation of motion of the $i$’th electron (2) in the scale-dependent dimensionless variables is reduced to
${\ddot{\rm\bf r}}_{i}^{*}+\,2\,u_{0}^{*}\,(1+\,u^{*}_{0}\,t^{*})^{-1}\,{\dot{\rm\bf r}}_{i}^{*}=\,(1+\,u^{*}_{0}\,t^{*})^{-3}\,{\rm\bf F}^{*}_{i}\,$, (10)
where dot denotes a derivative with respect to $t^{*}$. Therefore, expansion of the coordinate frame results in the appearance of an effective viscous force, proportional to the electron velocity, which is presented by the second term in the left-hand side of this equation. In fact, just this dissipative force is responsible for the electron deceleration and its subsequent trapping by the nearest ion.
The Coulomb forces, originally given by formula (3), after the normalization are reduced to
${\bf F}^{*}_{i}$ $=$ $\sum_{j=1}^{N}\,\frac{{\bf R}^{*}_{j}-{\bf r}^{*}_{i}}{|{\bf R}^{*}_{j}-{\bf r}^{*}_{i}|^{3}}\;+\,\sum_{j=1,\,j\neq i}^{N}\,\frac{{\bf r}^{*}_{i}-{\bf r}^{*}_{j}}{|{\bf r}^{*}_{i}-{\bf r}^{*}_{j}|^{3}}$ (11)
$+$ $\sum_{{\bf n}\,\in\,\mathbb{Z}^{3}\backslash\{{\bf 0}\}}\Bigg\{\sum_{j=1}^{N}\frac{({\bf R}^{*}_{j}+L^{*}\,{\bf n})-{\bf r}^{*}_{i})}{\big|({\bf R}^{*}_{j}+L^{*}\,{\bf n})-{\bf r}^{*}_{i}\big|^{3}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\>+\;\sum_{j=1}^{N}\frac{{\bf r}^{*}_{i}-({\bf r}^{*}_{j}+L^{*}\,{\bf n})}{\big|{\bf r}^{*}_{i}-({\bf r}^{*}_{j}+L^{*}\,{\bf n})\big|^{3}}\Bigg\}\>.$
As was already explained above, the sums over $\bf n$ are actually calculated over the cubic shells of increasing radii, composed of the “mirror” cells, until the required accuracy of convergence is achieved.
II.4 The method of integration
Following the standard procedure, the second-order differential equation (10) can be rewritten as a system of two first-order differential equations:
${\dot{\rm\bf r}}_{i}^{*}$ $\!\!\!\!=$ $\!\!\!\!{\rm\bf v}_{i}^{*},$ (12a)
${\dot{\rm\bf v}}_{i}^{*}$ $\!\!\!\!=$ $\!\!\!\!-2\,u_{0}^{*}s\,{\rm\bf v}_{i}^{*}+\,s^{3}\,{\rm\bf F}^{*}_{i},$ (12b)
where
$s(t^{*})=\,(1+\,u^{*}_{0}\,t^{*})^{-1}.$ (13)
Next, they are integrated numerically by the second-order Runge–Kutta method with a constant stepsize. In fact, this method is rather inefficient, and a much better computational performance could be achieved by using the higher-order integrators with the adaptive stepsize control. This was done, for example, in our previous works Dumin and Lukashenko (2022, 2024), where we employed the subroutines from Numerical Recipes library Press et al. (1992). Unfortunately, the adaptive stepsize control operates reliably only for the reflective boundaries and encounters serious problems in the case of periodic boundary conditions (for more details, see Appendix A in paper Dumin and Lukashenko (2022)). So, since the geometrical setup of our simulations (Fig. 2) inherently incorporates the periodic boundaries, we were enforced to use here a rather primitive integration scheme.
Yet another very subtle point in our modeling is usage of the exact (singular) Coulomb potentials of the charged particles, without any “softening” or cut-offs at the small distances, which was a common practice in many earlier papers Lankin and Norman (2009b); Niffenegger, Gilmore, and Robicheaux (2011); Tiwari, Shaffer, and Baalrud (2017). We preferred to avoid such a simplification, because formation of the bound states (resulting in the recombination) is rather sensitive to the specific form of the potential. Particularly, the closed trajectories are possible only in the exact Coulomb field Landau and Lifshitz (1976). (Yet another option for the formation of closed orbits is the potential of harmonic oscillator, proportional to $r^{2}$; but this is evidently beyond our physical context.)
Unfortunately, if the simulated plasma was initially cold enough, then the electrons began to fall onto the nearest ions, sometimes resulting in the almost head-on collisions, which could not be resolved by our integration algorithm. Such cases were identified by the sharp jumps in the kinetic energy of electrons and their asymmetric temporal profiles (with respect to the passage of pericenters); for the particular examples, see Appendix A in the present paper. Then, the integration was cancelled, and the corresponding initial conditions were discarded. So, we used only the “favorable” initial conditions, which did not lead to the failures. In fact, the number of failures was a few times greater than the number of successful attempts. Nevertheless, this did not increase appreciably the computational cost of our simulations, because the failures typically occurred at the very early stage of integration (about a few $\tau$ from the beginning), while we were interested in the long-term dynamics (several tens or hundreds $\tau$). So, a few failures per one successful attempt did not take too much time.
II.5 Initial conditions
An accurate specification of the initial conditions for ultracold plasmas is a rather nontrivial tack. For example, if the plasma is produced by instantaneous photoionization of neutral atoms, then the initial electron positions should be substantially correlated with the ionic ones, and distribution of the electron velocities will be non-Gaussian Niffenegger, Gilmore, and Robicheaux (2011). In the opposite case, when the ionization process takes some time, the released electrons will be mixed in space between the ions and approximately thermalized, resulting in the uniform coordinate and Gaussian velocity distributions. Since we are interested in the recombination events occurring after a quite long temporal evolution, exact specification of the initial conditions should not be important. So, for the sake of simplicity, we followed the second option. Namely, initial positions of the electrons $r^{*}_{i0}$ and ions $R^{*}_{i0}$ ($i=1,\dots,N$) were given by the random numbers distributed uniformly within the box, $L^{*}/2\leq(x^{*},y^{*},z^{*})\leq L^{*}/2$.
On the other hand, the initial electron velocities ${\dot{r}}^{*}_{i0}$ were specified by the normal (Gaussian) distribution with a root-mean-square deviation ${\sigma}_{v0}^{*}$ over each Cartesian coordinate. This quantity evidently characterizes the initial electron temperature. As was shown in our work Dumin (2011), the formal definition $T_{e}=(2/3)\,E_{\rm kin}/k_{\rm B}$ (where $E_{\rm kin}$ is the electron kinetic energy, and $k_{\rm B}$ is the Boltzmann constant) should work rather well even if the electrons strongly interact with ions. At last, as was already mentioned above, the ions are assumed to move by inertia due to their very large masses, i.e., in the co-moving reference frame they are at rest: $R^{*}_{i}\equiv\mbox{const}$ and ${\dot{R}}^{*}_{i}\equiv 0$.
III Results of the simulations
Figure 4: Temporal dependence of the kinetic (red curves) and potential (blue curves) energies of all particles in the basic cell at various “magnifications” (successive panels, from top to bottom). The regions of subsequent magnification are marked by the gray rectangles.
The simulations were performed for various numbers of the charged particles of each kind $N$ in the basic cell, from 10 to 1000. Since the employed algorithm of numerical integration without the adaptive stepsize control was computationally expensive, a sufficiently long interval of integration (about $1000\,\tau$) was achieved only for the small number of particles (e.g., $N=10$); and the most convincing patterns of recombination were obtained just in this case. On the other hand, when $N$ was substantially larger, the achievable interval became shorter by an order of magnitude; so that only the initial stage of recombination was observed, and it was difficult to derive a reliable conclusion on its total efficiency. (To avoid misunderstanding, let us remind once again that the actual number of particles taken into account in calculation of the Coulomb forces, including the “mirror” cells, was much greater: for example, even at $N=10$ it varied from a few hundred thousand to a few million, depending on the rate of convergence of the Coulomb sums at each particular step of integration.)
In the results presented below, we used the initial dispersion of the electron velocities ${\sigma}_{v0}^{*}=3.0$. In other words, the plasma was originally weakly non-ideal (${\Gamma}_{e}\sim 0.1$), since its initial kinetic energy exceeded the potential (Coulomb) energy only by one order of magnitude.
The overall plasma expansion rate was taken to be $u_{0}^{*}=0.1$. Then, “physical” size of the simulation cell (and, therefore, of the entire plasma cloud) changes in the course of time according to formula (7). Since the simulations presented below were conducted up to $t^{*}\approx 600$, the cloud expanded approximately by 60 times and its density, respectively, dropped by over 5 orders of magnitude. Such variation well corresponds to the typical experimental situations Killian et al. (1999, 2007).
Figure 5: Trajectory of the first (short-period) captured electron viewed in three coordinate planes at two different magnifications (three top and three bottom panels, respectively) as well as its kinetic energy (the fourth panels in the top and bottom). The region of formation of a captured state of the electron is marked by the red dotted circles.
Figure 6: Trajectory of the second (long-period) captured electron viewed in three coordinate planes (two top and left bottom panels) as well as its kinetic energy (right bottom panel). The region of formation of a captured state of the electron is marked by the red dotted circles.
Figure 4 shows a temporal behavior of the total kinetic and potential energies of all particles in the basic cell, including the energy of interaction with the “mirror” images. As can be expected, in the upper panel one can see initially a sharp decay in the kinetic energy, which is just an adiabatic cooling of the expanding plasma. As a result, the electron Coulomb coupling parameter ${\Gamma}_{e}$ quickly decays from ${\sim}0.1$ to unity and approximately stabilizes at this level.
The most interesting and nontrivial effect observed in this figure is the development of fast oscillations (a series of the sharp equidistant peaks), surviving up to the end of our simulation. Moreover, starting from $t^{*}\approx 250{-}300$ one can see onset of the second, much longer period of oscillations, which also survives up to the end of simulation. A presence of the corresponding two periods is clearly seen in the second, enlarged panel. At last, inspection of the third, most magnified panel gives us a hint that modulation of the kinetic and potential energies is caused just by the passages of an electron, captured into a strongly elliptical orbit, near the pericenter and apocenter of its trajectory.
A more careful analysis of the electron trajectories, presented in Figs. 5 and 6, confirms this conjecture. It is seen that the above-mentioned electrons—after some intervals of the “random walk”—really form the confined states. Their orbits are gradually contracting just because they are depicted in the co-moving reference frame, expanding with the plasma cloud. In fact, these pairs are detached from the overall expansion and have an approximately constant size; but they look contracting in the scalable coordinate system, where size of the basic cell $L^{*}$ remains constant.
The captured electrons should ultimately recombine with the corresponding ions into the neutral atoms when quantum effects will be taken into account. This corresponds to the well-known mechanism of the collisional two-stage recombination Massey and Burhop (1956); Smirnov (1981), whose total rate is governed mostly by the first stage (i.e., formation of the classical bound pairs). The trajectories of other eight electrons in the basic cell do not exhibit any features of localization. Consequently, we can estimate the total efficiency of recombination as 20%. This is actually the main result of our simulations.
A further inspection of the trajectories shows that the first electron (Fig. 5) became captured already at $t^{*}\approx 20{-}25$ (i.e., when the plasma cloud expanded by 3–3.5 times); but its orbit initially was very irregular and, therefore, did not exhibit a definite period. Besides, it is interesting that the long-period modulation of energy, caused by the second electron, is visible not only in the energy of the entire system (Fig. 4) but also in the energy of the first electron (bottom right panel in Fig. 5). In general, this is not surprising, because the quickly-oscillating first electron should naturally experience the long-period perturbations by the second electron.
At last, inspection of the bottom right panel in Fig. 6 shows that oscillations of energy of the second electron began to develop already at $t^{*}\approx 50{-}100$, while they are visible in the total energy (top panel in Fig. 4) only starting from $t^{*}\approx 250{-}300$. A gradual decay of the amplitude of these oscillations is evidently caused by a decreasing ellipticity of the orbit.
IV Discussion
In the present work we followed the well-known idea of collisional recombination Massey and Burhop (1956); Smirnov (1981) that the bound electron-ion pairs are formed initially due to the classical dynamics, and then almost all of them recombine to neutral atoms due to the quantum transitions (which are commonly assumed to occur much faster). So, the total efficiency of recombination should be equal approximately to the rate of formation of the classically-bound electron–ion pairs. (Of course, such an approach does not assume establishment of the thermodynamic equilibrium.)
The plasma in our simulations was initially weakly-coupled (${\Gamma}_{e}\sim 0.1$) but then quickly evolved to the moderately-coupled state (${\Gamma}_{e}\sim 1$), and the recombination processes developed just in this regime. Let us emphasize that the most of earlier UCP simulations were carried out in a box of fixed size for a set of “predefined” parameters (density, temperature, etc.). As distinct from that works, our method of “scalable frames” enables us to simulate the expanding plasmas by a self-consistent way, i.e., the recombination occurs at the time-dependent density, temperature, etc.; and all these parameters are determined “automatically” from the same simulation.
Although the sample of particles employed in our work was rather small, the results obtained turned out to be in agreement both with the experimental measurements Killian et al. (2001) as well as with the earlier numerical modeling Niffenegger, Gilmore, and Robicheaux (2011). In fact, the criterion of recombination in the last-cited paper was introduced rather arbitrarily as the percentage of electrons located closer than 0.2 Wigner–Seitz radius to the nearest ion. However, it is interesting that such choice of the “critical” radius enabled that authors to get approximately the same total efficiency of recombination (namely, $17{-}18\,\%$) as in our ab initio simulation.
Yet another frequently-used criterion Lankin and Norman (2009a) is to find the electrons that performed a specified minimal number of revolutions (e.g., 1, 2, 4, or 6) about the nearest ion and to assume that they will ultimately recombine to the neutral atoms. In fact, this approach is even more popular and often considered as justified better. Unfortunately, analysis of the results of our simulations did not provide a clear support in favor of such criterion. So, this subject still needs to be studied in more detail.
As was already mentioned above, the employed numerical algorithm is time-consuming: even at the relatively small number of particles in the basic cell, each computational run on the ordinary PC (without parallelization) took a few months. So, when the number of particles was much greater, e.g. 100 or 1000, we were able to perform simulations only for the considerably shorter time intervals, about $100\,\tau$, when the recombination processes were not complete yet. The low speed of calculations was caused by the fact that we used neither the adaptive stepsize control for integration, nor Ewald summation technique for the Coulomb sums. However, one can expect that it will be possible in future to incorporate these methods into our algorithm of the scalable (co-moving) reference frames, which should make the such simulations much faster.
In fact, it is not sufficiently clear how sensitive the recombination statistics is to the accurate treatment of the long-range forces and if the corresponding procedures can be reasonably simplified. On the one hand, the bound states cannot be evidently formed by the two-body interactions (since the total energy of the pair will be conserved in this case). So, the multi-particle interactions should be important in any case Massey and Burhop (1956). On the other hand, one can try to describe them in a simpler approximation, for example, by using the “minimum-image convention”, i.e., summation over the particles located within the distance $\pm L/2$ from the given particle. To the best of our knowledge, the earlier attempts by other authors to employ the simplified summation schemes did not enable them to get the sufficiently stable electron–ion pairs, but this subject still needs to be studied in more detail.
Besides, it should be mentioned that a purely classical electron–ion Coulomb system is not stable in the genuine thermodynamic limit, because the energy is unbounded from below and the trajectories can collapse; so that the quantum-mechanical effects will be essential for the stability of matter. However, this argumentation is irrelevant to the case under consideration: our system cannot reach the thermodynamic equilibrium because it steadily expands up to infinity. In such a case, the most evident stabilizing factor—preventing collapse of the electron trajectories to the singularity—is an approximate conservation (or, more exactly, adiabatic invariance) of the angular momentum $M$ of a captured electron with respect to the corresponding ion, which enters into the effective potential Landau and Lifshitz (1976):
$U_{\rm eff}=\,\frac{M^{2}}{2m_{e}r^{2}}\,-\,\frac{e^{2}}{r}\,.$ (14)
Then, the first term in the right-hand side prevents the electron orbit from the collapse as long as $M$ remains the adiabatic invariant. In other words, the mechanism of (quasi-)stabilization in this case is substantially different from any kinds of quantum corrections, such as a finite electron thermal de Broglie wavelength Demyanov and Levashov (2025), etc.
Let us emphasize that the numerical failures during “head-on” collisions, which were discarded in our simulations, are absolutely irrelevant to the above-mentioned collapse of the classical Coulomb systems in the thermodynamic limit. Really, the discarded runs typically involved the "numerical pumping" of energy (i.e., ejection of the electrons) rather than a collapse of their trajectories to the singularity of the potential. In fact, we have already encountered the same numerical failures in our previous works about the clusterized plasmas Dumin and Lukashenko (2022, 2024). This problem was easily resolved there by using the integration with the adaptive stepsize control (ASSC), particularly, the subroutines from the Numerical Recipes library Press et al. (1992). Unfortunately, these subroutines become unreliable in the case of periodic boundary conditions, which are essential for the present modeling. So, as a simplest remedy, we just discarded the unfavorable initial conditions. In principle, it should be possible to modify the ASSC algorithms for the periodic boundary conditions, but this will require a lot of work on the development of new software.
In fact, a lot of current works employ a much simpler approach, based on the truncated Coulomb interactions; and they successfully reproduce many properties of the ultracold plasmas. However, we were afraid to do so in simulation of the recombination, because the truncated Coulomb potential results in the inherently unclosed classical orbits and, in principle, this can affect the subsequent quantum transitions to the ground atomic states.
V Conclusions
  1. 1.
    We suggested a new approach for modeling of the expanding ultracold plasmas, which is based on the “scalable” reference frames, where the effect of expansion formally manifests itself as an effective viscous force in the co-moving coordinates. This enabled us to simulate a non-stationary behavior of such plasmas by a self-consistent way, when all physical parameters (density, temperature, etc.) are determined in the course of time by the same model rather than specified in advance for a set of simulations in the fixed-size boxes.
  2. 2.
    To the best of our knowledge, the present work is the first successful ab initio simulation of recombination in the expanding ultracold plasmas, which clearly demonstrates how some of electrons become localized and form the stable compact pairs with the respective ions. Such pairs are assumed to recombine subsequently to the ground atomic states due to the quantum transitions. It is important that we did not employ here any artificial criteria for the formation of classical electron–ion pairs (such as a minimal number of revolutions or a maximal distance between the particles).
  3. 3.
    The efficiency of recombination at the realistic experimental conditions obtained by our method turned out to be in agreement both with laboratory measurements and the earlier semi-empirical estimates. For example, if the initial value of the electron coupling parameter ${\Gamma}_{e}$ was about 0.1 or somewhat less, and the plasma cloud expanded by 50–60 times in radius (and, correspondingly, its density dropped by over 5 orders of magnitude), about 20% of the original electrons recombined, which coincides perfectly with the experiment Killian et al. (2001). The earlier numerical modeling Niffenegger, Gilmore, and Robicheaux (2011) resulted in a similar value, 17–18%; but this coincidence should be taken with caution, because the corresponding simulation was carried out in a box of fixed size, and the recombined particles were identified by the semi-empirical criterion.
  4. 4.
    Unfortunately, the proposed modeling scheme is rather expensive computationally: for example, a single run takes a few months on the ordinary PC. However, we expect that its efficiency can be improved in future with an optimized software, which should employ the Ewald summation technique for the Coulomb forces and the adaptive stepsize control for integration.
Acknowledgements.
YVD is grateful to A.A. Bobrov, P.R. Levashov, S.A. Mayorov, J.-M. Rost, S.A. Saakyan, U. Saalmann, V.S. Vorob’ev, B.B. Zelener, and B.V. Zelener for fruitful discussions and valuable suggestions as well as to O.E. Stakheev for the assistance in numerical calculations.
Author declarations
Conflict of interest
The authors have no conflicts to disclose.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Examples of the “favorable” and “unfavorable” initial conditions
Figure 7: Temporal variations of the kinetic and potential energies of all particles in the basic cell in the case of “favorable” initial conditions (which are the same as in Figs. 46).
Figure 8: Temporal variations of the kinetic and potential energies in the case of “unfavorable” initial conditions when integration was performed with standard (top panel) and enhanced accuracy (bottom panel).
As was already mentioned in Sec. II.4, the employed integrator was sometimes unable to resolve reliably the close passages of electrons near the ions in the very beginning of simulations, and the respective initial conditions were discarded. Here we provide the corresponding examples.
Figure 7 represents the case of a well-resolved passage of an electron near the pericenter of its strongly-elliptic orbit at the sufficiently large time, when the stable electron–ion pairs were already formed. (The initial conditions are the same as in Figs. 46.) One can see here two major features in variations of the kinetic and potential energies: Firstly, the energetic peaks are symmetric, as should be expected for motion in the elliptic orbit (if weak perturbations from the remote particles are ignored). Secondly, the peaks in kinetic and absolute value of potential energies are approximately of the same magnitude, as should be naturally expected from conservation of the total energy. Really, the potential (Coulomb) energy is actually negative and, in the case of strongly-elliptic orbit, the total energy of the pair is much less than the above-mentioned peak values (i.e., roughly speaking, is close to zero).
An absolutely different situation can be seen in Fig. 8, which corresponds to the case of “unfavorable” initial conditions. In general, a very chaotic pattern of variations is not surprising, because this is an early stage of the plasma cloud expansion, when relaxation did not complete yet and the plasma remained rather hot. One can see in the top panel a sharp asymmetric jump of the kinetic energy at $t^{*}=10.7$, which is evidently unphysical and caused just by numerical errors. Really, this increase in the kinetic energy (about 20 arbitrary units) is an order of magnitude greater than the associated increase in the absolute value of potential energy (about 2 units), which violates the energy conservation condition. Next, the peak of potential energy is approximately symmetric (because position of the electron always changes smoothly with respect to the ion), but the kinetic energy exhibits a strongly asymmetric jump (step). In fact, such cases are well known in UCP simulations and commonly called the “numerical heating”.
However, when the accuracy of numerical integration is substantially improved (bottom panel in Fig. 8), the above-mentioned jump transforms into a tiny symmetric peak, whose amplitude is comparable to the peak of potential energy, as should be expected for a close interparticle collision. Therefore, the problem was purely numerical. Unfortunately, because of the limited computational resources, we were unable to perform the entire set of simulations with the substantially improved accuracy. So, the “unfavorable” cases were just disregarded.